OpenStudy (anonymous):

cos(3x) = -1 Part 1: using an appropriate inverse trigonometric expression, write an equation that defines the value of 3x. Part 2: solve this equation to find all possible values of the angle 3x. Part 3: use algebra to find all values of x between 0 and 2pi that satisfy this equation.

zepdrix (zepdrix):

Remember inverse functions back from algebra? :) To find the inverse we would `swap` the variables, and then solve for our new y. Same idea here, we change our cosine to inverse cosine, and swamp the argument with the other side. \[\large \cos(3x)=-1 \qquad\to\qquad \arccos(-1)=3x\]

zepdrix (zepdrix):

swap* not swamp lol

OpenStudy (anonymous):

That makes a lot of sense, Thank you! :D Could you help me with the other questions too?

zepdrix (zepdrix):

Hmm part 2, let's seeee. I actually think it's easier to leave it in cosine form to solve the second part. We have \(\large \cos(3x)=-1\) Where 3x is just some angle, let's call it theta for right now,\[\large \cos \theta=-1\] Do you remember your special angles? What angle when you take the cosine of it, gives us -1?

zepdrix (zepdrix):

Oh actually, that's part 3.. my bad. For part 2, they just want you to solve for x. So in this case, we would just divide both sides by 3 right? \[\large x=\frac{\arccos(-1)}{3}\]

OpenStudy (anonymous):

so x= pi/3 ?

zepdrix (zepdrix):

Yesss very good! :)

OpenStudy (anonymous):

Thank you! So for part 3, is the special angle pi?

zepdrix (zepdrix):

Yes it is. If we had done it that way, we would have determined that \(\large \theta=\pi\). Or in other words, \(\large 3x=\pi\). And then dividing by 3 would get us the same answer. But it looks like you already figured that out based on the other method we used from here\[\large x=\frac{\arccos(-1)}{3}\] So it's all good :D

OpenStudy (anonymous):

Okay! Thanks so much! I totally understand what to do now! :D